X squared + y squared = 4
X - 2y = 1
I'm really confused with this whole problem

Answer by ankor@dixie-net.com(22740)

X squared + y squared = 4
X - 2y = 1
:
Two equations
x^2 + y^2 = 4
and
x - 2y = 1
Add 2y to both sides
x = 2y + 1
Now you can substitute (2y+1) for x in the 1st equation
(2y+1)^2 + y^2 = 4
FOIL (2y+1)(2y+1)
4y^2 + 4y + 1 + y^2 = 4
combine like terms to form a quadratic equation
4y^2 + y^2 + 4y + 1 - 4 = 0
5y^2 + 4y - 3 = 0
Use the quadratic formula to find y; a=5; b=4; c=-3;I got:
y = -1.272
and
y = +.472
:
Find x using both solutions
when y = -1.272
x = 2(-1.272) + 1
x = -2.544 + 1
x = -1.544
:
when y = .472
x = 2(.472) + 1
x = 1.944
:
:
You should check both pairs of solutions in the 1st equations, will be close but not exact. we had to round off the solutions